OpenStudy (dls):

What is the condition for a body projected from the surface of the earth to behave as a black hole?

OpenStudy (dls):

@Mashy

OpenStudy (anonymous):

:O .. WHAT?!??.. what has.. black hole gotta do with PROJECTION? black hole purely depends on the density!.. you need some enormous density for it to be considered as a blackhole

OpenStudy (dls):

yeah i want the answer in terms of radius/density only :o

OpenStudy (anonymous):

there is something called as schwarzchild radius.. google it :P

OpenStudy (dls):

oho..mast..if we project it greater than equal to speed of light so it might behave as black hole..how do we do it numerically?

OpenStudy (anonymous):

i think you need to use relavistic speeds and stuff.. don't tell me this is a part of your syllabus? :O

OpenStudy (dls):

\[\LARGE V_e > c\]

OpenStudy (dls):

hmm it is :P

OpenStudy (dls):

\[\LARGE \sqrt{\frac{2GM}{R}}\ge c\]

OpenStudy (anonymous):

OMg now i understand what you are trying to say.. your question was so poorly constructed man :-/ The escape velocity > C.. blah!!

OpenStudy (dls):

that is the condition for it to behave as a black hole?just asking..thats not the Q

OpenStudy (anonymous):

Yes yes.. that is the expression for escape velocity from the surface of a planet or star or any mass for that matter :P which has a mass M and radius R..

OpenStudy (anonymous):

if the escape velocity is greater than light, then it means its a black hole

OpenStudy (dls):

how can a body become black hole because of its speed :S and can u find the value of radius of the body ?

OpenStudy (anonymous):

a body can become a black hole because of its speed Cause as you move your mass increases :P.. but thats relativity.. ! Please post exactly what is the question?? the question you stated seems absurd!

OpenStudy (dls):

Q is correct :/ and mass will increase to? black hole o.O why will it absorb

OpenStudy (anonymous):

when you say condition? what exactly is it that you are asking for? :O and mass increase is given by this relation \[m(moving) = m(rest) /\sqrt{1- v ^{2}/c^{2}}\]

OpenStudy (dls):

wtf

OpenStudy (anonymous):

I told you.. thats the expression you get from Einstien's theory of relativity!